OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −2

OpenStudy (anonymous):

It wants you to use substitution so lets get the value of y and substitute it into the other equation 2x + y = 1 4x + 2y = −2 2x + y = 1 y = -2x + 1 Now that we know what y is lets use it as a substitution for y in the other equation 4x + 2(-2x+1) = −2 Can you solve for x now?